OpenStudy (anonymous):

standard form of (-1,-7)(1,-1)??

jimthompson5910 (jim_thompson5910):

First find the slope m = (y2-y1)/(x2-x1) m = (-1-(-7))/(1-(-1)) m = (-1+7)/(1+1) m = 6/2 m = 3 So the slope of the line that passes through the two points is 3 Now use the slope along with the first point to find the equtaion of the line in slope intercept form y = mx+b -7 = (3)(-1)+b -7 = -3+b -7+3 = b -4 = b b = -4 So the equation of the line in slope intercept form is y = 3x-4 Now convert to standard form y = 3x-4 -3x + y = -4 3x - y = 4